OpenStudy (senia01):

Check my answers? A line has a slope of 7 and contains the point (5, −9). Which equations represent the line? Choose all answers that are correct. A. −7x + y = −44 <---- B. 7x – y = 44 <---- C. (y − 9) = 7(x + 5) D. (y + 9) = 7(x − 5)

OpenStudy (mathstudent55):

\(y = mx + b\) m = 7, so \(y = 7x + b\) Using point (5, -9) \(-9 = 7(5) + b\) \(-9 = 35 + b\) \(b = -44\) The equation of the line is \(y = 7x - 44\)

OpenStudy (texaschic101):

-7x + y = -44 y = 7x - 44 (5,-9)....x = 5 and y = -9 -9 = 7(5) - 44 -9 = 35 - 44 -9 = -9 (correct)....this is correct y + 9 = 7(x - 5) y + 9 = 7x - 35 y = 7x - 35 - 9 y = 7x - 44.....this is one also

OpenStudy (mathstudent55):

A. −7x + y = −44 y = -7x - 44 <----- correct B. 7x – y = 44 y = 7x - 44 <------ correct C. (y − 9) = 7(x + 5) y - 9 = 7x + 35 y = 7x + 44 <----- incorrect D. (y + 9) = 7(x − 5) y + 9 = 7x - 35 y = 7x - 44 <------ correct Choices A, B, and D are the answer.

OpenStudy (senia01):

So c is the only wrong one

OpenStudy (mathstudent55):

Exactly.

OpenStudy (texaschic101):

yep

OpenStudy (senia01):

Thank you (again)

OpenStudy (mathstudent55):

yw